# NarrativeBridge: Enhancing Video Captioning with Causal-Temporal Narrative

### v1

@@ 3.2.1 Stage 1: Causal Visual Encoding
We instantiate $E_{cause}$ and $E_{effect}$ as two separate instances of the CLIP-ViT (ViT-B/32) model [37], adapted from CLIP4Clip [29]. For both Cause and Effect text encoders ($T_{cause}$ and $T_{effect}$ respectively), we employ a 12-layer transformer with a hidden size of $512$ and $8$ attention heads, using the weights from the pre-trained CLIP [37] text encoder. As in CLIP [37], the [EOS] token’s representation from the last layer of the text encoders is used as the feature representation for the input text. Similar to CLIP4Clip, we adopt cosine similarity to measure the similarity between the video embedding (mean pooled) and the text embedding. Given a video embedding $\hat{r}_{i}$ and a text embedding $t_{j}$, the similarity function is defined as $s(\hat{r}_{i},t_{j})=\frac{t_{j}^{\top}\hat{r}_{i}}{|t_{j}||\hat{r}_{i}|}$. During training, each model is optimized using a contrastive loss, as in CLIP [37], over its respective cause/effect portion from CTN captions. For a batch of $N$ video-cause text pairs $(v_{i},c_{i})$ and video-effect text pairs $(v_{i},e_{i})$, the contrastive losses for the cause and effect video encoders are defined as:
@@ 3.2.2 Stage 2: CTN Caption Generation
In the second stage (see Figure 3), we employ a sequence-to-sequence transformer-based network, consisting of two separate encoders ($Enc_{cause}$ and $Enc_{effect}$) for processing cause and effect video features, respectively, and a decoder $Dec$ for generating the final CTN captions. The encoded cause and effect video representations are concatenated before being sent to the decoder. Both encoders and the decoder are initialized with the weights from the pre-trained Uni-VL model [28]. Given a video $v$, we first extract the cause and effect video features using the pre-trained $E_{cause}$ and $E_{effect}$ encoders: $F_{cause}=E_{cause}(v),$ (5) $F_{effect}=E_{effect}(v).$ (6)
@@ 4 Experiments
We conduct extensive experiments to validate our proposed CEN architecture and the new CTN video captioning benchmark. This section details the datasets used, the experimental setup, and the evaluation metrics. We also present comprehensive results and evaluation against the state-of-the-art video captioning methods demonstrating the performance of our network on the CTN benchmark.
@@ 4.1 Datasets
We evaluate our CTN captions and CEN on two widely-used video captioning datasets: 1) MSR-VTT [55], consists of 10,000 videos with 200,000 human-annotated captions, and MSVD [7], with 1,970 videos focused on human activities with 78,800 captions. The captions in these datasets lack explicit representations of the causal relationships, temporal dynamics, motivations, intentions, and the narrative of events and human actions in the videos. To address these limitations, we generate CTN captions using the Mixtral of Experts LLM [20] as detailed in Section 3.1.
@@ 4.2 Evaluation Metrics
For quantifying the performance of our method CEN against the SOTA methods, we use three metrics as per [6]; CIDEr (C) [47]: Measures alignment of generated captions with references by highlighting frequently occurring terms, capturing the ability to reproduce salient causal-temporal narrative elements. ROUGE-L (R-L) [23]: Evaluates the longest common subsequence between generated and reference captions, considering both precision and recall for assessing semantic similarity, including temporal dynamics and causal relationships. SPICE (S) [3]: Evaluates semantic quality by considering the overlap of scene graphs between generated and reference captions, effectively assessing the ability to capture and express underlying causal relationships and event sequences.
@@ 4.3 Implementation Details
We generate the CTN captions using the Mixtral of Experts LLM [20], running on A100-80GB GPUs. All models, including our proposed CEN architecture and the SOTA models (SEM-POS [32], AKGNN [17], and GIT [51]), are implemented using PyTorch. For the SOTA models, we follow the hyperparameter settings specified in their respective methods for training on the MSR-VTT and MSVD datasets. Our CEN model is trained using the Adam optimizer with learning rates of $1\times 10^{-4}$ (stage 1) and $1\times 10^{-6}$ (stage 2) and a batch size of 64 for 10 epochs (stage 1) and 50 epochs (stage 2). Further details are provided in the Appendix A.
@@ 4.4 Results and Discussion
We comprehensively evaluate the performance of our proposed CEN architecture against several SOTA methods, including SEM-POS [32], AKGNN [17], and GIT [51], on our CTN captions generated for MSR-VTT [55] and MSVD [7] benchmark datasets.
Quantitative Results: CEN outperformed all SOTA methods across all metrics and datasets as shown in Table 2. On MSVD, CEN surpassed the next best (GIT) by 3.95 ROUGE-L, 17.88 CIDEr, and 3.67 SPICE points. On MSR-VTT, CEN led GIT by 3.39 ROUGE-L, 17.44 CIDEr, and 2.06 SPICE points. These significant gains highlight CEN’s effectiveness in capturing causal narratives and temporal dynamics. To gain further insights into the effectiveness of the CTN captions in capturing causal-temporal narratives, we visualize the video feature representations learned by CLIP-ViT encoders trained on CTN and Original captions using UMAP [31] dimensionality reduction from high dimension onto a 2D plane. In Figure 4(a), we compare the representations learned from CTN captions (red) and Original captions (blue) on the MSR-VTT dataset. The non-overlapping feature spaces indicate that CTN captions capture the causal and temporal relationships not present in the original captions. In Figure 4(b), we visualize the representations learned from the cause (black) and effect (orange) parts of the CTN captions. The near-complete overlap suggests a strong correlation between the cause and effect components, aligning with the inherent structure of causal-temporal narratives and supporting the design of the CEN architecture.
Table 1: Comparison of our CEN architecture against SOTA methods on the MSVD and MSR-VTT datasets. The best results in each category are in bold. R-L, C, and S denote ROUGE-L, CIDEr, and SPICE scores, respectively. Method MSVD MSRVTT R-L ()($\uparrow$)( ↑ ) C ()($\uparrow$)( ↑ ) S ()($\uparrow$)( ↑ ) R-L ()($\uparrow$)( ↑ ) C ()($\uparrow$)( ↑ ) S ()($\uparrow$)( ↑ ) SEM-POS [32] 25.39 37.16 14.46 20.11 26.01 12.09 AKGNN [17] 25.11 35.08 14.55 21.42 25.90 11.99 GIT [51] 27.51 45.63 15.58 24.51 32.43 13.70 CEN (Ours) 31.46 63.51 19.25 27.90 49.87 15.76 Table 2: Ablation study results on the MSVD and MSR-VTT datasets. $\text{E}_{\text{JointCE}}$, Only $\text{E}_{\text{cause}}$ and Only $\text{E}_{\text{effect}}$ are variants of our CEN architecture, while Zero Shot v.v. and Fine-tune v.v. represent cross-dataset evaluation settings. The best results in each category are in bold. Method MSVD MSRVTT R-L ()($\uparrow$)( ↑ ) C ()($\uparrow$)( ↑ ) S ()($\uparrow$)( ↑ ) R-L ()($\uparrow$)( ↑ ) C ()($\uparrow$)( ↑ ) S ()($\uparrow$)( ↑ ) $\text{E}_{\text{JointCE}}$ 30.93 55.72 17.04 27.34 45.97 15.07 Only $\text{E}_{\text{cause}}$ 30.72 56.42 18.43 27.19 47.10 15.21 Only $\text{E}_{\text{effect}}$ 30.70 57.14 17.89 27.24 45.58 15.19 Zero Shot v.v. 27.16 39.65 14.45 24.73 29.76 12.26 Fine-tune v.v. 31.78 65.60 19.39 27.47 47.74 15.65 CEN (Ours) 31.46 63.51 19.25 27.90 49.87 15.76
@@
Qualitative Analysis: Qualitative examples, in Figure 5, demonstrate CEN’s strength in accurately articulating causal relationships and event sequences, while SOTA methods struggle. For instance, in (a), CEN captures the fatality move’s causal effect, unlike baselines. In (c), CEN links the kick’s precision to scoring, which others miss. Overall, quantitative and qualitative results showcase CEN’s superior performance in understanding and generating causal-temporal narrative video captions compared to existing SOTA methods. More results are provided in Appendix F.
@@
Ablation Results: To examine the effectiveness of CEN, we compare it with 4 baselines in Table 2:
@@ 6 Conclusion
In this paper, we introduced for the first time a Causal-Temporal Narrative (CTN) captions benchmark and proposed a novel Cause-Effect Network (CEN) tailored for causal-temporal narrative video captioning. This work finds vast applications in automated video summarization, question- answering, assistive technologies, and surveillance. The CTN captions benchmark provides a comprehensive testbed for evaluating video understanding models’ ability to grasp complex temporal and causal dynamics, which will be released for research purpose on https://narrativebridge.github.io/. And CEN explicitly models cause-effect relationships and temporal dynamics to generate rich, contextually relevant descriptions capturing nuanced causal-temporal narrative in videos, demonstrating significant performance improvement over SOTA methods (Section 4). NarrativeBridge lays the foundation for a paradigm shift, where models comprehend underlying causal-temporal narrative driving events, unlocking new frontiers in contextually aware human-machine interactions with video.
For future work, we aim to integrate CTN caption generation with existing image captioning techniques, to annotate unlabeled videos with causal-temporal narrative labels. A few frames of the video will be labelled using off-the-shelf image captioning methods, and CTN caption generation will exploit the labelled frames to generate one coherent caption for the unlabelled video (see Appendix E). This synergistic approach opens new avenues for comprehensive video understanding and annotation, enabling more robust and accurate video analysis pipelines.
7 Acknowledgment
@@
Appendix
Appendix A Implementation Details
We first extract frames from the video clip and then encode them using the Cause and Effect Video Encoders. We adopt a sampling strategy where we sample one frame per second, with a maximum of 20 frames. This approach ensures that the sampled frames cover the entire duration of the video, providing a comprehensive representation of the video content while maintaining computational efficiency. For the fine-tuning experiment in Table 2, all the hyperparameters remain the same except the learning rate i.e. 0.0000005. For our experimentation of the Cause and Effect Network (CEN), we use standard splits of the MSVD [7] and MSRVTT [55] datasets. For MSVD, we use 1200 videos for training, 100 for validation, and 670 for testing. For MSRVTT, we use 6513 videos for training, 497 for validation, and 2990 for testing. These splits are commonly used in video captioning research to ensure fair comparison with other methods. All the experiments in Stage 1 and Stage 2 of our CEN are run using A100-80GB and RTX 3090-24GB GPUs respectively. We implement SEM-POS [32], AKGNN [17] and GIT [51] using RTX 3090-24GB, A100-80GB and A100-80GB GPUs respectively.
Appendix B Prompt Design Process
@@
Appendix C Comparison of Large Language Models for CTN Caption Generation
To evaluate the effectiveness of large language models (LLMs) in generating high-quality Causal-Temporal Narrative (CTN) captions, we compare the performance of two open source state-of-the-art LLMs at the time of experimentation: Mixtral of Experts [20], which we utilized in our CTN caption generation pipeline, and Llama2-70b [46].

### v2

@@ 3.2.1 Stage 1: Causal Visual Encoding
We instantiate $E_{cause}$ and $E_{effect}$ as two separate instances of the CLIP-ViT (ViT-B/32) model Radford et al. (2021), adapted from CLIP4Clip Luo et al. (2022) (see Appendix A.6 for more details). For both Cause and Effect text encoders ($T_{cause}$ and $T_{effect}$ respectively), we employ a 12-layer transformer with a hidden size of $512$ and $8$ attention heads, using the weights from the pre-trained CLIP Radford et al. (2021) text encoder. As in CLIP Radford et al. (2021), the [EOS] token’s representation from the last layer of the text encoders is used as the feature representation for the input text. Similar to CLIP4Clip, we adopt cosine similarity to measure the similarity between the video embedding (mean pooled) and the text embedding. Given a video embedding $\hat{r}_{i}$ and a text embedding $t_{j}$, the similarity function is defined as $s(\hat{r}_{i},t_{j})=\frac{t_{j}^{\top}\hat{r}_{i}}{|t_{j}||\hat{r}_{i}|}$. During training, each model is optimized using a contrastive loss, as in CLIP Radford et al. (2021), over its respective cause/effect portion from CTN captions. For a batch of $N$ video-cause text pairs $(v_{i},c_{i})$ and video-effect text pairs $(v_{i},e_{i})$, the contrastive losses for the cause and effect video encoders are defined as:
@@ 3.2.2 Stage 2: CTN Caption Generation
In the second stage (see Figure 3), we employ a sequence-to-sequence transformer-based network, consisting of two separate encoders ($Enc_{cause}$ and $Enc_{effect}$) for processing cause and effect video features, respectively, and a decoder $Dec$ for generating the final CTN captions. The encoded cause and effect video representations are concatenated before being sent to the decoder. Both encoders and the decoder are initialized with the weights from the pre-trained Uni-VL model Luo et al. (2020). Given a video $v$, we first extract the cause and effect video features using the pre-trained $E_{cause}$ and $E_{effect}$ encoders:
$F_{cause}=E_{cause}(v),$ (5)
$F_{effect}=E_{effect}(v).$ (6)
@@ 4 Experiments
We conduct extensive experiments to validate our proposed CEN architecture and the new CTN video captioning benchmark. This section details the datasets used, the experimental setup, and the evaluation metrics. We also present comprehensive results and evaluation against the state-of-the-art video captioning methods and vision-language models (VLMs) demonstrating the performance of our network on the CTN benchmark.
@@ 4.1 Datasets
We evaluate our CTN captions and CEN on two widely-used video captioning datasets: 1) MSR-VTT Xu et al. (2016), consists of 10,000 videos with 200,000 human-annotated captions, and MSVD Chen & Dolan (2011), with 1,970 videos focused on human activities with 78,800 captions. The captions in these datasets lack explicit representations of the causal relationships, and temporal dynamics which we address by generating CTN captions as detailed in Section 3.1.
@@ 4.2 Evaluation Metrics
For quantifying the performance of our method CEN against the SOTA methods, we use three metrics as per Celikyilmaz et al. (2020); CIDEr (C) Vedantam et al. (2015): Measures alignment of generated captions with references by highlighting frequently occurring terms, capturing the ability to reproduce salient causal-temporal narrative elements. ROUGE-L (R-L) Lin (2004): Evaluates the longest common subsequence between generated and reference captions, considering both precision and recall for assessing semantic similarity, including temporal dynamics and causal relationships. SPICE (S) Anderson et al. (2016): Evaluates semantic quality by considering the overlap of scene graphs between generated and reference captions, effectively assessing the ability to capture and express underlying causal relationships and event sequences.
@@ 4.3 Implementation Details
We generate the CTN captions using the Mixtral of Experts LLM Jiang et al. (2024), running on A100-80GB GPUs. All models, including our proposed CEN architecture, the SOTA models (SEM-POS Nadeem et al. (2023), AKGNN Hendria et al. (2023), and GIT Wang et al. (2022)) and the VLMs VideoLLaVA Lin et al. (2023) and ShareGPT4Video Chen et al. (2024) are implemented using PyTorch. For the SOTA models, we follow the hyperparameter settings specified in their respective methods for training on the MSR-VTT and MSVD datasets. Our CEN model is trained using the Adam optimizer with learning rates of $1\times 10^{-4}$ (stage 1) and $1\times 10^{-6}$ (stage 2) and a batch size of 64 for 10 epochs (stage 1) and 50 epochs (stage 2). For comparison with recent Vision-Language Models (VLMs), we fine-tune VideoLLaVA Lin et al. (2023) and ShareGPT4Video Chen et al. (2024) using both LoRA and simple fine-tuning approaches on our CTN benchmark. We use the recommended hyperparameters for each model during fine-tuning. Further details are provided in the Appendix A.1.
@@ 4.4 Results and Discussion
We comprehensively evaluate the performance of our proposed CEN architecture against several SOTA methods, including SEM-POS Nadeem et al. (2023), AKGNN Hendria et al. (2023), and GIT Wang et al. (2022) and the VLMs VideoLLaVA Lin et al. (2023) and ShareGPT4Video Chen et al. (2024) on our generated CTN captions benchmark datasets.
Table 1: Comparison of our CEN architecture against SOTA methods and VLMs on the MSVD and MSR-VTT datasets. The best results in each category are in bold. R-L, C, and S denote ROUGE-L, CIDEr, and SPICE scores, respectively. Method MSVD MSRVTT R-L ()($\uparrow$)( ↑ ) C ()($\uparrow$)( ↑ ) S ()($\uparrow$)( ↑ ) R-L ()($\uparrow$)( ↑ ) C ()($\uparrow$)( ↑ ) S ()($\uparrow$)( ↑ ) SEM-POS 25.39 37.16 14.46 20.11 26.01 12.09 AKGNN 25.11 35.08 14.55 21.42 25.90 11.99 GIT 27.51 45.63 15.58 24.51 32.43 13.70 VideoLLaVA (Zero-shot) 21.80 30.55 14.67 19.33 16.24 12.49 VideoLLaVA (LoRA FT) 24.56 34.98 15.41 21.21 18.97 13.28 VideoLLaVA (Simple FT) 25.61 36.12 16.09 22.18 19.98 13.07 ShareGPT4Video (Zero-shot) 21.66 27.06 14.06 20.27 17.08 12.21 ShareGPT4Video (LoRA FT) 24.39 30.72 14.83 22.09 19.83 13.02 ShareGPT4Video (Simple FT) 25.32 31.67 14.92 23.01 20.76 13.28 CEN (Ours) 31.46 63.51 19.25 27.90 49.87 15.76 Table 2: Ablation study results on the MSVD and MSR-VTT datasets. $\text{E}_{\text{JointCE}}$, Only $\text{E}_{\text{cause}}$ and Only $\text{E}_{\text{effect}}$ are variants of our CEN architecture, while Zero Shot v.v. and Fine-tune v.v. represent cross-dataset evaluation settings. The best results in each category are in bold. Method MSVD MSRVTT R-L ()($\uparrow$)( ↑ ) C ()($\uparrow$)( ↑ ) S ()($\uparrow$)( ↑ ) R-L ()($\uparrow$)( ↑ ) C ()($\uparrow$)( ↑ ) S ()($\uparrow$)( ↑ ) $\text{E}_{\text{JointCE}}$ 30.93 55.72 17.04 27.34 45.97 15.07 Only $\text{E}_{\text{cause}}$ 30.72 56.42 18.43 27.19 47.10 15.21 Only $\text{E}_{\text{effect}}$ 30.70 57.14 17.89 27.24 45.58 15.19 Zero Shot v.v. 27.16 39.65 14.45 24.73 29.76 12.26 Fine-tune v.v. 31.78 65.60 19.39 27.47 47.74 15.65 CEN (Ours) 31.46 63.51 19.25 27.90 49.87 15.76
@@
4.4.1 Quantitative Results
CEN outperformed all SOTA methods and the VLMs across all metrics and datasets as shown in Table 2. On MSVD, CEN surpassed the next best (GIT) by 3.95 ROUGE-L, 17.88 CIDEr, and 3.67 SPICE points. On MSR-VTT, CEN led GIT by 3.39 ROUGE-L, 17.44 CIDEr, and 2.06 SPICE points. These significant gains highlight CEN’s effectiveness in capturing causal narratives and temporal dynamics.
We compare CEN’s performance against zero-shot and fine-tuned (FT) versions of VideoLLaVA and ShareGPT4Video. CEN consistently outperforms these models, even after fine-tuning, demonstrating the effectiveness of our specialized architecture for causal-temporal narrative generation. For example, on MSR-VTT, CEN achieves a CIDEr score of 49.87, compared to 19.98 for VideoLLaVA (Simple FT) and 20.76 for ShareGPT4Video (Simple FT).
To demonstrate generalization capability, we perform cross-dataset evaluations (Table 2). CEN shows strong zero-shot performance (Zero Shot v.v.) and significant improvements after fine-tuning (Fine-tune v.v.), outperforming models trained directly on the target datasets. For instance, when trained on MSR-VTT and tested on MSVD, CEN achieves a CIDEr score of 65.60 after fine-tuning, compared to 45.63 for GIT trained directly on MSVD.
To gain further insights into the effectiveness of the CTN captions in capturing causal-temporal narratives, we visualize the video feature representations learned by CLIP-ViT encoders trained on CTN and Original captions using UMAP McInnes et al. (2018) dimensionality reduction from high dimension onto a 2D plane. In Figure 4(a), we compare the representations learned from CTN captions (red) and Original captions (blue) on the MSR-VTT dataset. The non-overlapping feature spaces indicate that CTN captions capture the causal and temporal relationships not present in the original captions. In Figure 4(b), we visualize the representations learned from the cause (black) and effect (orange) parts of the CTN captions. The near-complete overlap suggests a strong correlation between the cause and effect components, aligning with the inherent structure of causal-temporal narratives and supporting the design of the CEN architecture.
@@ 4.4.1 Quantitative Results
4.4.2 Qualitative Results
Qualitative examples, in Figure 5, demonstrate CEN’s strength in accurately articulating causal relationships and event sequences, while SOTA methods and recent VLMs struggle. For example: (a) Video game scene: CEN accurately captures the causality ("performs a fatality move" causing "killing another"), while other methods miss this crucial relationship. SEM-POS and GIT provide overly simplistic descriptions, while AKGNN generates irrelevant details. VLMs (Video-LLaVA and ShareGPT4Video) misinterpret the scene entirely.
(b) Paper folding: CEN correctly identifies both the cause (folding paper) and effect (creating a paper airplane). Other methods either miss the causal relationship or misinterpret the action (e.g., ShareGPT4Video’s "small square" instead of an airplane).
(c) Soccer scene: CEN accurately links the precise kick to the goal, capturing both cause and effect. Other methods either focus on a single action or provide generic descriptions of a soccer game.
(d) Singing performance: CEN captures the boy’s decision to perform (cause) and the audience’s reaction (effect). Other methods mostly describe the act of singing without the causal-temporal context.
Overall, quantitative and qualitative results showcase CEN’s superior performance in understanding and generating causal-temporal narrative video captions compared to existing SOTA methods and recent VLMs. More results are provided in Appendix A.9.
4.4.3 Ablation Results
To examine the effectiveness of CEN, we compare it with 4 baselines in Table 2:
@@ 6 Conclusion
In this paper, we introduced for the first time a Causal-Temporal Narrative (CTN) captions benchmark and proposed a novel Cause-Effect Network (CEN) tailored for causal-temporal narrative video captioning. This work finds vast applications in automated video summarization, question-answering, assistive technologies, surveillance, and educational content creation. The CTN captions benchmark provides a comprehensive testbed for evaluating video understanding models’ ability to grasp complex temporal and causal dynamics, which will be released for research purpose on https://narrativebridge.github.io/. And CEN explicitly models cause-effect relationships and temporal dynamics to generate rich, contextually relevant descriptions capturing nuanced causal-temporal narrative in videos, demonstrating significant performance improvement over SOTA methods (Section 4). NarrativeBridge lays the foundation for a paradigm shift, where models comprehend underlying causal-temporal narrative driving events, unlocking new frontiers in contextually aware human-machine interactions with video.
For future work, we aim to integrate CTN caption generation with existing image captioning techniques, to annotate unlabeled videos with causal-temporal narrative labels. A few frames of the video will be labelled using off-the-shelf image captioning methods, and CTN caption generation will exploit the labelled frames to generate one coherent caption for the unlabelled video (see Appendix A.8). This synergistic approach opens new avenues for comprehensive video understanding and annotation, enabling more robust and accurate video analysis pipelines.
Acknowledgement
@@
Appendix A Appendix
A.1 Implementation Details
We first extract frames from the video clip and then encode them using the Cause and Effect Video Encoders. We adopt a sampling strategy where we sample one frame per second, with a maximum of 20 frames. This approach ensures that the sampled frames cover the entire duration of the video, providing a comprehensive representation of the video content while maintaining computational efficiency. For the fine-tuning experiment in Table 2, all the hyperparameters remain the same except the learning rate i.e. 0.0000005. For our experimentation of the Cause and Effect Network (CEN), we use standard splits of the MSVD Chen & Dolan (2011) and MSRVTT Xu et al. (2016) datasets. For MSVD, we use 1200 videos for training, 100 for validation, and 670 for testing. For MSRVTT, we use 6513 videos for training, 497 for validation, and 2990 for testing. These splits are commonly used in video captioning research to ensure fair comparison with other methods. All the experiments in Stage 1 and Stage 2 of our CEN are run using A100-80GB and RTX 3090-24GB GPUs respectively. We implement SEM-POS Nadeem et al. (2023), AKGNN Hendria et al. (2023) and GIT Wang et al. (2022) using RTX 3090-24GB, A100-80GB and A100-80GB GPUs respectively.
For comparison with Vision-Language Models (VLMs), we implement two fine-tuning approaches: LoRA Fine-Tuning and Simple Fine-Tuning using A100-80GB GPUs. LoRA Fine-Tuning is applied specifically to the LLM component, with a learning rate of 2e-4 for LoRA parameters. Simple Fine-Tuning is applied to the entire model, using an AdamW Loshchilov (2017) optimizer with a cosine learning rate schedule (initial learning rate: 1e-3, warmup ratio: 0.03).
A.2 Prompt Design Process
@@
A.3 Comparison of Large Language Models for CTN Caption Generation
To evaluate the effectiveness of large language models (LLMs) in generating high-quality Causal-Temporal Narrative (CTN) captions, we compare the performance of two open source state-of-the-art LLMs at the time of experimentation: Mixtral of Experts Jiang et al. (2024), which we utilized in our CTN caption generation pipeline, and Llama2-70b Touvron et al. (2023).